# Maximum gap in complex Ginibre matrices

### v1

@@ 1. Introduction
In the literature on extremal eigenvalue gaps, the majority of works have concentrated on models with one-dimensional spectrum. The classical Gaussian random matrix ensembles, which include the Gaussian Orthogonal Ensemble (GOE), Gaussian Unitary Ensemble (GUE), and Gaussian Symplectic Ensemble (GSE), are paradigmatic examples. The correlation functions for these ensembles admit explicit formulas, allowing their eigenvalue behavior to be studied through direct computations. As a result, the distributions of the smallest gaps are known for each of the classical Gaussian ensembles [49, 7, 21, 20]. The problem of determining the distributions of the largest gaps seems more challenging. Currently, these distributions are known only for the GUE; they were determined in [22], following earlier work in [7] that obtained the asymptotic size of the largest gaps. Further, it was shown in [10] that these results on the largest and smallest gaps are universal in the sense that they hold for all Wigner matrices, assuming the entry distributions posses a density satisfying a weak smoothness condition. [^1] Roughly speaking, Wigner matrices are symmetric random matrices whose upper-triangular entries are independent with mean zero and variance one. In the case of the largest gap, universality without the smoothness condition was shown in [34]; this result encompasses, for example, Bernoulli random matrices.
@@ 1.1. Proof ideas
We consider an auxiliary point process, denoted $\Xi_{c,n}$ below, that consists of the set of $n$ GinUE eigenvalues thinned to retain only those points whose near-neighbor spacing is greater than a certain distance $r_{n,c}$, defined precisely below, where $c>0$ is a parameter introduced for technical reasons. Our main result, Theorem 1, states that on any set $B\subset\mathbb{C}$ such that $\sup_{z\in B}|z|<1$, the Kantorovich–Rubinstein distance between $\Xi_{c,n}$ and a suitably chosen Poisson process $\zeta_{c}$ tends to zero as $n\rightarrow\infty$. Routine arguments then allow one to deduce the leading-order asymptotic for the nearest-neighbor spacing in $B$, stated in Corollary 2.
Our approach is similar to [44], which determined the leading-order asymptotic for the largest gap for the limiting process of the GinUE eigenvalues. However, [44] focused on deriving a general Poisson process approximation result for stabilizing functionals of determinantal point processes. The error terms in this approximation are relatively straightforward to bound for the limiting point process by exploiting its stationarity. Our emphasis is different. We rely on the simpler Poisson approximation result [9, Theorem 3.1], but face new technical challenges because the eigenvalue process of the GinUE is not stationary. In particular, we would like to apply [9, Theorem 3.1] to $\Xi_{c,n}$, but that result requires a coupling of $\Xi_{c,n}$ to its Palm version $\Xi^{z}_{c,n}$ for every $z\in B$. Due to the non-stationarity of the eigenvalue process, we can only couple these processes after restricting them to a set that omits a small ball around $z$. By adapting the proof of [9, Theorem 3.1], we show that the small set where we cannot construct a coupling can be neglected, then bound the error terms arising on the complementary large set (without relying on stationarity, in contrast to [44]). These bounds rely on large deviations estimates for the probability that a given region contains at most two GinUE eigenvalues.
@@ 1.2. Related works
The circular ensembles, which have one-dimensional spectra, are another class of classical matrix ensembles. The distribution of the smallest gaps for the CUE was derived in [7], along with the order of the largest gaps. The distributions of the largest gaps were found in [22]. Small gaps of the circular $\beta$-ensembles for integer $\beta$, a more general model which specializes to the CUE in the case $\beta=2$, were studied in [23]. The largest and smallest gaps for a multi-matrix ensemble that generalizes the GUE were studied in [24].
In models where obtaining the full distribution of the smallest gap is challenging, one can instead ask for high-probability lower bounds. These have been derived for symmetric random matrices in [42, 36] and matrices of independent entries in [25, 38]. For the Coulomb gas model, which generalizes the GinUE, various separation estimates for the particles have been shown in [3, 4, 48, 5]. Largest gap behavior is closely related to existence of large zero-free regions, which have been studied in [14, 2, 1, 13].
@@ 2. Model and main results
$\rho^{(m)}_{\nu}(z_{1},\dots,x_{m})=\det(K(z_{i},z_{j}))_{i,j=1}^{m},\quad z_{1},\dots,z_{m}\in\mathbb{C},\,\quad m\in\mathbb{N},$
It is well known (see, e.g., [8, Theorem 4.3.10]) that $\xi_{n}$ is a determinantal point process with kernel
$K_{n}(z,w)=\sum_{k=1}^{n}\varphi_{k}(z)\overline{\varphi_{k}(w)},\quad\text{where }\quad\varphi_{k}(z):=\frac{1}{\sqrt{\pi(k-1)!}}z^{k-1}e^{-\frac{1}{2}|z|^{2}}.$
In this article, we study the largest gap in the process $\xi_{n}$. To do so, we consider the point process-valued function
(1) $\Xi_{c,n}[\omega]:=\sum_{z\in\omega\cap B_{\sqrt{n}}(0)}\mathds{1}\{(\omega\setminus\{z\})(B_{r_{c,n}(z)}(z))=0\}\,\delta_{z/\sqrt{n}},\quad c>0,\quad\omega\in\mathbf{N},$
and let $\Xi_{c,n}:=\Xi_{c,n}[\xi_{n}]$. The sequence $(r_{c,n}(z))_{n\in\mathbb{N}}$ is given by
(2) $r_{c,n}(z):=\inf\{r>0:\,n\mathbb{P}(\xi_{n}^{z!}(B_{r}(z))=0)\rho_{n}(z)\leq c\},\quad z\in\mathbb{C}.$
Here $\rho_{n}(z):=K_{n}(z,z)$ is the one-point correlation function for $\xi_{n}$ and $\xi_{n}^{z!}$ denotes the reduced Palm process of $\xi_{n}$ at $z$ (see Section 3 for its definition). [^2] The infimum in the definition of $r_{c,n}$ is achieved by some $r>0$. This follows from the fact, noted below, that $\xi_{n}^{z!}$ is a determinantal point process with kernel (8). Then the function
$\mathbf{L}_{c}(A):=\int_{\sqrt{n}A}\mathbb{P}(\xi_{n}^{z!}(B_{r_{c,n}}(z))=0)\rho_{n}(z)\,\mathrm{d}z=c|A|,\quad A\subset B_{1}(0),$
where $|A|$ denotes Lebesgue measure of $A$, does not depend on $n$. We will see in the next section that $\mathbf{L}_{c}$ is the intensity measure for $\Xi_{c,n}$ for all $n\in\mathbb{N}$.
@@
where Lip is the class of all measurable 1-Lipschitz functions $h:\widehat{\mathbf{N}}\to\mathbb{R}$ with respect to the total variation between measures $\omega_{1},\omega_{2}$ on $\mathbb{C}$, which is given by
@@
We now state our two results, which are proved in Section 5. The first shows that $\Xi_{c,n}$ may be approximated by a Poisson process at a resolution small enough to capture the behavior of the largest eigenvalue gap.
@@
Fix $c>0$ and recall that $\xi_{n}$ denotes the point process of eigenvalues for the $n$-dimensional complex Ginibre matrix, and that $\Xi_{c,n}:=\Xi_{c,n}[\xi_{n}]$ is the process defined at (1). Then for all $B\subset B_{1}(0)$ such that $\sup_{z\in B}|z|<1$, there exists $C(B)>0$ such that for all $\varepsilon>0$ and $n\in\mathbb{N}$,
(4) ${d_{\mathrm{KR}}}(\Xi_{c,n}\cap B,\zeta_{c}\cap B)\leq Cn^{\varepsilon-1/16},$
where $\zeta_{c}$ is a Poisson process with intensity measure $\mathbf{L}_{c}$. Further, we have for all $c>0$ that as $n\to\infty$,
(5) $\sup_{z\in\sqrt{n}B}\Big{|}\frac{r_{c,n}(z)^{4}}{4\log n}-1\Big{|}\to 0.$
Our second result is a consequence of the previous theorem. It obtains the asymptotic scaling of the largest distance from a point $x\in\xi_{n}$ to its nearest neighbor. This scaling is different than the one for independent points, which is order $\log n$ [15].
@@
$\frac{1}{4\log n}\max_{z\in\xi_{n}\cap\sqrt{n}B}\min_{w\in\xi_{n}\setminus\{x\}}|z-w|^{4}\stackrel{{\scriptstyle\mathbb{P}}}{{\longrightarrow}}1.$
@@ 3. Preliminaries
3.1. Palm calculus and negative association
Since all arguments are analogous for each choice of $c>0$, we fix $c$ and drop the subscript $c$ in the following. Moreover, we write $B_{r_{n}}(z)$ instead of $B_{r_{n}(z)}(z)$.
We introduce two Palm processes that will play fundamental roles in the proof of Theorem 1. By [32, Chapter 6] there exist point processes $\xi_{n}^{z}$ and $\xi_{n}^{z,\varnothing}$ for all $z\in\mathbb{C}$ such that for all $f:\mathbb{C}\times\mathbf{N}\to[0,\infty)$,
(6) $\mathbb{E}\Big{[}\int_{\sqrt{n}A}f(z,\xi_{n})\xi_{n}(\mathrm{d}z)\Big{]}$ $=\int_{\sqrt{n}A}\mathbb{E}[f(z,\xi_{n}^{z})]\rho_{n}(z)\,\mathrm{d}z,$
[…]
$=\int_{A}\mathbb{E}[f(z,\xi_{n}^{\sqrt{n}z,\varnothing})]\mathbf{L}(\mathrm{d}z).$
To get (7), we first used that [32, Lemma 6.2(ii)] implies $\delta_{z}\in\xi_{n}^{z}$ almost surely for all $n\in\mathbb{N}$, allowing us to define the reduced Palm process $\xi_{n}^{z!}:=\xi_{n}^{z}-\delta_{z}$ and hence $\mathbf{L}(\mathrm{d}z)$. Then, in the second equality in (7), we used that $\mathbf{L}(\mathrm{d}z)$ is the intensity measure for $\Xi_{c,n}$, which follows from (6) by choosing
$f(z,\xi_{n})=\mathds{1}\{(\xi_{n}-\delta_{z})(B_{r_{n}}(z))=0\}.$
Again using [32, Lemma 6.2(ii)], we have $\delta_{z}\in\xi_{n}^{z,\varnothing}$ for all $n$ and can define $\xi_{n}^{z!,\varnothing}:=\xi_{n}^{z,\varnothing}-\delta_{z}$. In the following, we write $P_{n}$, $P_{n}^{z!}$ and $P_{n}^{z!,\varnothing}$ for the distributions of $\xi_{n}$, $\xi_{n}^{z!}$ and $\xi_{n}^{z!}$, respectively.
@@ 3.1. Palm calculus and negative association
(see [47, Theorem 1.7]). By [28, Theorem 3] (see also [40]), the process $\xi_{n}^{x!}$ is stochastically dominated by $\xi_{n}$ (denoted by $P_{n}^{x!}\leq P_{n}$), which means that

### v2

@@ 1. Introduction
In the literature on extremal eigenvalue gaps, the majority of works have concentrated on models with one-dimensional spectrum. The classical Gaussian random matrix ensembles, which include the Gaussian Orthogonal Ensemble (GOE), Gaussian Unitary Ensemble (GUE), and Gaussian Symplectic Ensemble (GSE), are paradigmatic examples. The correlation functions for these ensembles admit explicit formulas, allowing their eigenvalue behavior to be studied through direct computations. As a result, the distributions of the smallest gaps are known for each of the classical Gaussian ensembles [52, 8, 22, 21]. The problem of determining the distributions of the largest gaps seems more challenging. Currently, these distributions are known only for the GUE; they were determined in [23], following earlier work in [8] that obtained the asymptotic size of the largest gaps. Further, it was shown in [11] that these results on the largest and smallest gaps are universal in the sense that they hold for all Wigner matrices, assuming the entry distributions posses a density satisfying a weak smoothness condition. [^1] Roughly speaking, Wigner matrices are symmetric random matrices whose upper-triangular entries are independent with mean zero and variance one. In the case of the largest gap, universality without the smoothness condition was shown in [36]; this result encompasses, for example, Bernoulli random matrices. Very recently, universality for the smallest gap was extended to a large class of matrices containing non-smooth distributions, include those with sufficiently many atoms [53].
@@ 1.1. Proof ideas
We consider an auxiliary point process, denoted $\Xi_{n}$ below, that consists of the set of $n$ GinUE eigenvalues thinned to retain only those points whose near-neighbor spacing is greater than a certain distance $r_{n}$, defined precisely below. Our main result, Theorem 1, states that on any set $B\subset\mathbb{C}$ such that $\sup_{z\in B}|z|<1$, the Kantorovich–Rubinstein distance between $\Xi_{n}$ and a suitably chosen Poisson process $\zeta_{c}$ tends to zero as $n\rightarrow\infty$. Routine arguments then allow one to deduce the leading-order asymptotic for the nearest-neighbor spacing in $B$, stated in Corollary 2.
Our approach is similar to [47], which determined the leading-order asymptotic for the largest gap for the limiting process of the GinUE eigenvalues. However, [47] focused on deriving a general Poisson process approximation result for stabilizing functionals of determinantal point processes. The error terms in this approximation are relatively straightforward to bound for the limiting point process by exploiting its stationarity. Our emphasis is different. We rely on the simpler Poisson approximation result [10, Theorem 3.1], but face new technical challenges because the eigenvalue process of the GinUE is not stationary. In particular, we would like to apply [10, Theorem 3.1] to $\Xi_{n}$, but that result requires to control the symmetric difference of $\Xi_{n}$ and its reduced Palm version at every $z\in B$, which seems difficult to accomplish. However, due to the stabilization property of the eigenvalue process, we can construct a coupling (with a certain monotonicity property) of these processes after restricting them to a set that omits a small ball around $z$. By adapting the proof of [10, Theorem 3.1], we show that the small set where we cannot construct a coupling can be neglected, then bound the error terms arising on the complementary large set (without relying on stationarity, in contrast to [47]). These bounds rely on large deviations estimates for the probability that a given region contains at most two GinUE eigenvalues.
@@ 1.2. Related works
The circular ensembles, which have one-dimensional spectra, are another class of classical matrix ensembles. The distribution of the smallest gaps for the CUE was derived in [8], along with the order of the largest gaps. The distributions of the largest gaps were found in [23]. Small gaps of the circular $\beta$-ensembles for integer $\beta$, a more general model which specializes to the CUE in the case $\beta=2$, were studied in [24]. The largest and smallest gaps for a multi-matrix ensemble that generalizes the GUE were studied in [25]. For the spherical ensemble, a determinantal process of points on the sphere that can be realized as the stereographic projection of the eigenvalues of a certain random matrix, the asymptotic area of the largest empty cap was derived in [3].
In models where obtaining the full distribution of the smallest gap is challenging, one can instead ask for high-probability lower bounds. These have been derived for symmetric random matrices in [45, 39] and matrices of independent entries in [26, 41]. For the Coulomb gas model, which generalizes the GinUE, various separation estimates for the particles have been shown in [4, 5, 51, 6], and recently the convergence of the smallest gaps process to a Poisson process was established in [15], generalizing the GinUE result of [49]. Largest gap behavior is closely related to existence of large zero-free regions, which have been studied in [14, 2, 1, 13].
@@ 2. Model and main results
$\rho^{(m)}_{\nu}(z_{1},\dots,x_{m})=\det(K(z_{i},z_{j}))_{i,j=1}^{m},\quad z_{1},\dots,z_{m}\in\mathbb{C},\,\quad m\in\mathbb{N}.$
It is well known (see, e.g., [9, Theorem 4.3.10]) that $\xi_{n}$ is a determinantal point process with kernel
(1) $K_{n}(z,w)=\sum_{k=1}^{n}\varphi_{k}(z)\overline{\varphi_{k}(w)},\quad\text{where }\quad\varphi_{k}(z):=\frac{1}{\sqrt{\pi(k-1)!}}z^{k-1}e^{-\frac{1}{2}|z|^{2}}.$
In this article, we study the largest gap in the process $\xi_{n}$. To do so, we let $\kappa>1$ be a parameter and consider the point process-valued function
(2) $\Xi_{n}[\omega]:=\sum_{z\in\omega\cap B_{\sqrt{n}}(0)}\mathds{1}\{(\omega\setminus\{z\})(B_{r_{n}(z)}(z))=0\}\,\delta_{z/\sqrt{n}},\quad\omega\in\mathbf{N},$
where the sequence $(r_{n}(z))_{n\in\mathbb{N}}$ is given by
(3) $r_{n}(z):=\inf\{r>0:\,n\mathbb{P}(\xi_{n}^{z!}(B_{r}(z))=0)\rho_{n}(z)\leq\kappa\},\quad z\in\mathbb{C}.$
Here $\rho_{n}(z):=K_{n}(z,z)$ is the one-point correlation function for $\xi_{n}$, and $\xi_{n}^{z!}$ denotes the reduced Palm process of $\xi_{n}$ at $z$ (see Section 3 for its definition). We define $\Xi_{n}=\Xi_{n}[\xi_{n}]$ and let $L_{n}(A)=\mathbb{E}[\Xi_{n}(A)]$ denote the intensity measure of $\Xi_{n}$, where $A\subset\mathbb{C}$ is assumed to be measurable. While $\Xi_{n}$, $r_{n}$, and $L_{n}$ depend on $\kappa$, we omit this dependence in the notation.
@@
where Lip is the class of all measurable 1-Lipschitz functions $h\colon\widehat{\mathbf{N}}\to\mathbb{R}$ with respect to the total variation between measures $\omega_{1},\omega_{2}$ on $\mathbb{C}$, which is given by
@@
We now state our two results, which are proved in Section 5. The first shows that $\Xi_{n}$ may be approximated by a Poisson process at a resolution small enough to capture the behavior of the largest eigenvalue gap.
@@
Fix $s\in(0,1)$, $\epsilon>0$, and $\kappa>1$, and recall that $\xi_{n}$ denotes the point process of eigenvalues for the $n$-dimensional complex Ginibre matrix, and that $\Xi_{n}:=\Xi_{n}[\xi_{n}]$ is the process defined at (2). There exists $C(s,\epsilon,\kappa)>1$ such that for all $B\subset B_{1}(0)$ satisfying $\sup_{z\in B}|z|<s$, we have for all $n\in\mathbb{N}$ that
(5) ${d_{\mathrm{KR}}}(\Xi_{n}\cap B,\zeta_{\kappa}\cap B)\leq Cn^{\epsilon-1/16},$
where $\zeta_{\kappa}$ is a Poisson process with intensity measure $L_{n}$. Further, we have for all $\kappa>1$ that as $n\to\infty$,
(6) $\sup_{z\in\sqrt{n}B}\Big|\frac{r_{n}(z)^{4}}{4\log n}-1\Big|\to 0.$
Our second result is a consequence of the previous theorem. It obtains the asymptotic scaling of the largest distance from a particle in $\xi_{n}$ to its nearest neighbor. This scaling is different than the one for independent points, which is order $\log n$ [16].
@@
$\frac{1}{4\log n}\max_{z\in\xi_{n}\cap\sqrt{n}B}\min_{w\in\xi_{n}\setminus\{z\}}|z-w|^{4}\stackrel{{\scriptstyle\mathbb{P}}}{{\longrightarrow}}1.$
@@ 3. Preliminaries
3.1. Palm calculus
We begin by introducing three point processes that will play fundamental roles in the proof of Theorem 1.
Definition 3.
Let $z\in\mathbb{C}$. The following equalities are assumed to hold for all measurable $f\colon\mathbb{C}\times\mathbf{N}\to[0,\infty)$.
  • (a)
    $\xi_{n}^{z}$ is a Palm process of $\xi_{n}$ at $z$ if it satisfies
    (7) $\mathbb{E}\left[\int_{B_{\sqrt{n}}(0)}f(z,\xi_{n})\xi_{n}(\mathrm{d}z)\right]$ $=\int_{B_{\sqrt{n}}(0)}\mathbb{E}\big[f(z,\xi_{n}^{z})\big]\rho_{n}(z)\,\mathrm{d}z.$
  • (b)
    $\Xi_{n}^{z}$ is a Palm process of $\Xi_{n}$ at $z$ if it satisfies
    (8) $\mathbb{E}\left[\int_{B_{1}(0)}f(z,\Xi_{n})\Xi_{n}(\mathrm{d}z)\right]$ $=\int_{B_{1}(0)}\mathbb{E}\big[f(z,\Xi_{n}^{z})\big]L_{n}(\mathrm{d}z).$
  • (c)
    $\xi_{n}^{z,\varnothing}$ is a Palm process of $\xi_{n}$ with respect to $\Xi_{n}$ at $z$ if it satisfies [^2] The superscripted $\varnothing$ serves as a reminder of the indicator function for the event $(\xi_{n}\setminus\{z\})(B_{r_{n}}(z))=\varnothing$ present in the definition of $\Xi_{n}$.
    (9) $\mathbb{E}\left[\int_{B_{1}(0)}f(z,\xi_{n})\Xi_{n}(\mathrm{d}z)\right]$ $=\int_{B_{1}(0)}\mathbb{E}\big[f(z,\xi_{n}^{z,\varnothing})\big]L_{n}(\mathrm{d}z).$
The processes $\xi_{n}$, $\xi_{n}^{z!}:=\xi_{n}^{z}\setminus\{z\}$, $\Xi_{n}^{z!}:=\Xi_{n}^{z}\setminus\{z\}$ and $\xi_{n}^{z!,\varnothing}:=\xi_{n}^{z,\varnothing}\setminus\{z\}$ are called reduced Palm processes.
It follows from standard results that for all $z\in\mathbb{C}$ and $n\in\mathbb{N}$, there exist point processes $\xi_{n}^{z}$, $\Xi_{n}^{z}$, and $\xi_{n}^{z,\varnothing}$ satisfying parts (a), (b), and (c) of Definition 3, respectively [33, Chapter 6]. Further, [33, Lemma 6.2(ii)] implies $z\in\xi_{n}^{z}$ almost surely, and parallel claims hold for $\Xi_{n}^{z}$ and $\xi_{n}^{z,\varnothing}$.
Remark 4.
Comparing (8) and (9), we find that that $\Xi_{n}[\xi_{n}^{z,\varnothing}]$ is a Palm process of $\Xi_{n}$ at $z$. Indeed, if we choose $\tilde{f}(x,\omega):=f(x,\Xi_{n}[\omega])$ in (9), then its left-hand side coincides with the left-hand side of (8), and comparing their right-hand sides gives the claim.
We write $P_{n}$, $P_{n}^{z}$ and $P_{n}^{z,\varnothing}$ for the distributions of $\xi_{n}$, $\xi_{n}^{z}$ and $\xi_{n}^{z,\varnothing}$, respectively, and $P_{n}^{z!}$ and $P_{n}^{z!,\varnothing}$ for the distributions of the associated reduced Palm processes. For brevity, we write $B_{r_{n}}(z)$ instead of $B_{r_{n}(z)}(z)$.
@@ 3.1. Palm calculus
(see [50, Theorem 1.7]).
Lemma 5.
For every $s\in(0,1)$ and $\kappa>1$, there exists $C(s,\kappa)>1$ such that for all $z\in B_{s\sqrt{n}}(0)$ and $n\geq C$,
(11) $n\mathbb{P}(\xi_{n}^{z!}(B_{r_{n}}(z))=0)\rho_{n}(z)=\kappa.$
Proof.
It is straightforward to see that $\rho_{n}(z_{n})\uparrow\frac{1}{\pi}$ uniformly as $n\to\infty$ for all sequences $(z_{n})_{n=1}^{\infty}$ such that $z_{n}\in\sqrt{n}B$. Then there exists $C(s)>1$ such that for all $z\in B_{s\sqrt{n}}(0)$ and $n\geq C$,
(12) $\frac{1}{2\pi}\leq\rho_{n}(z)\leq\frac{1}{\pi}.$
Using that $\xi^{z!}_{n}$ is determinantal (see (10)), we have for every $r>0$ that
(13) $\mathbb{P}(\xi^{z!}_{n}(B_{r}(z))=0)=\frac{1}{n!}\int_{B_{r}(z)^{c}}\dots\int_{B_{r}(z)^{c}}\det(K^{z}_{n}(z_{i},z_{j}))_{i,j=1}^{m}\,dz_{1}\dots dz_{m}.$
Together with (10), this representation shows that the function $g(r)=\mathbb{P}(\xi^{z!}_{n}(B_{r}(z))=0)$ is a continuous function of $r$ for $r\in[0,\infty)$. Further, it is clear from (13) that $g(0)=1$ and $\lim_{r\rightarrow\infty}g(r)=0$. The claim follows from combining these limits with (12). ∎
Lemma 6.
For every $s\in(0,1)$ and $\kappa>1$, there exists $C(s,\kappa)>1$ such that for all $A\subset B_{s}(0)$ and $n\geq C$, it holds that $L_{n}(A)=c|A|$.
Proof.
By the definition of $\Xi_{n}$ and (7) with $f(z,\omega)=\mathds{1}\{\omega(B_{r_{n}}(z))=1\}$ we obtain
$L_{n}(A)$ $=\mathbb{E}\Big[\sum_{z\in\xi_{n}}\mathds{1}\{z/\sqrt{n}\in A\}\mathds{1}\{\xi_{n}(B_{r_{n}}(z))=1\}\Big]$
$=\mathbb{E}\Big[\int_{\sqrt{n}A}\mathds{1}\{\xi_{n}(B_{r_{n}}(z))=1\}\,\xi_{n}(\mathrm{d}z)\Big]=\int_{\sqrt{n}A}\mathbb{P}(\xi_{n}^{z}(B_{r_{n}}(z))=1)\rho_{n}(z)\,\mathrm{d}z.$
Now the assertion follows from (3) and Lemma 5. ∎
3.2. Stochastic domination
By [29, Theorem 3] (see also [43]), the process $\xi_{n}^{x!}$ is stochastically dominated by $\xi_{n}$ (denoted by $P_{n}^{x!}\leq P_{n}$), which means that